# AnyEdit: Edit Any Knowledge Encoded in Language Models

### v2

@@ A.1 Datasets
UnKEBench (Deng et al., 2024) constructs a dataset containing 1,000 counterfactual unstructured texts, where knowledge is presented in an unstructured and relatively lengthy form, going beyond simple knowledge triplets or linear fact chains. These texts originate from ConflictQA (Xie et al., 2023), a benchmark specifically designed to distinguish LLMs’ parameter memory from anti-memory. This approach is crucial for preventing the model from merging knowledge obtained during pretraining with knowledge acquired during the editing process. Moreover, it addresses the key challenge of determining whether the model learns target knowledge during training or editing, ensuring a clear boundary between pretraining knowledge and edited knowledge.
@@ A.2 Evaluation Metrics
$\mathbb{E}\left\{y=\mathop{\arg\max}\limits_{y^{\prime}}\mathbb{P}_{f}(y^{\prime}\left|x\right.)\right\}.$ (10)
@@
$\mathbb{E}\left\{y=\mathop{\arg\max}\limits_{y^{\prime}}\mathbb{P}_{f}(y^{\prime}\left|x_{e}\right.)\right\}.$ (11)
@@
$\mathbb{E}\left\{y_{loc}=\mathop{\arg\max}\limits_{y^{\prime}}\mathbb{P}_{f}(y^{\prime}\left|x_{loc}\right.)\right\}.$ (12)
@@ A.3 Baseline Methods
  • AlphaEdit (Fang et al., 2024) is a method designed to mitigate interference in LLM lifelong knowledge editing. It introduces a null-space projection mechanism that ensures parameter updates preserve previously edited knowledge while incorporating new updates. AlphaEdit has demonstrated state-of-the-art (SOTA) performance across multiple evaluation metrics while maintaining strong transferability. We follow the original paper’s hyperparameter configuration in our implementation.
  • UnKE (Deng et al., 2024) improves knowledge editing by refining both the layer and token dimensions. In the layer dimension, it replaces local key-value storage with a non-local block-based mechanism, enhancing the representation capability of key-value pairs while integrating attention-layer knowledge. In the token dimension, it replaces ”term-driven optimization” with ”cause-driven optimization,” which directly edits the final token while preserving contextual coherence. This eliminates the need for explicit term localization and prevents context loss.
@@ A.4 Implementation Details
Our AnyEdit and AnyEdit* primarily follow the baseline configurations of MEMIT and UnKE, while other baselines adhere to their original implementation settings. All experiments were conducted on a single A100 GPU (80GB).
@@ B.1 Locate-Then-Edit Paradigm
Following prior works on model editing, the detailed descriptions of specific methods in this section are based on MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024) and ECE (Zhang et al., ). We adhere to their formulations and methodological explanations to ensure consistency and clarity in presenting these approaches.
@@
Methods such as ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), and AlphaEdit (Fang et al., 2024) focus on triplets $(s,r,o)$, selecting the last token of the subject $s$ as the target token. In contrast, UnKE (Deng et al., 2024) extends to unstructured text, using the last token of $x_{e}$ as the target.
@@ C.1.2 Case 2
MEMIT To find the value of $log_{y}(x^{4})$, we can use the change of base formula and the given information. Given: $log_{x}(y^{3})=2$. Using the change of base formula, we can rewrite this as: $log_{x}(y^{3})=(log_{y}(y^{3}))/(log_{y}(x))$. Since $log_{y}(y^{3})=3$ (because $y^{3}$ is the argument raised to the power of 3), we have: $2=3/(log_{y}(x))$. Now, we can solve for $log_{y}(x)$: $log_{y}(x)=3/2$. Next, we want to find $log_{y}(x^{4})$. Using the power rule of logarithms, we can write: $log_{y}(x^{4})=4*log_{y}(x)$. Substituting the value we found for $log_{y}(x)$: $log_{y}(x^{4})=4*(3/2)$. Simplifying: $log_{y}(x^{4})=6$. Therefore, the value of $log_{y}(x^{4})$ is $6$.
AlphaEdit To find the value of $log_{y}(x^{4})$, we can use the change of base formula and the given information: $log_{x}(y^{3})=2$. Using the change of base formula, we can rewrite this as: $y^{3}=x^{2}$. Now, we want to find $log_{y}(x^{4})$. Using the properties of logarithms, we can rewrite this as: $log_{y}(x^{4})=4*log_{y}(x)$. We can express $log_{y}(x)$ in terms of $log_{x}(y)$ using the change of base formula: $log_{y}(x)=1/log_{x}(y)$. Since we know that $log_{x}(y^{3})=2$, we can write: $log_{x}(y^{3})=3*log_{x}(y)=2$. Therefore: $log_{x}(y)=2/3$. Substituting this into our expression for $log_{y}(x)$: $log_{y}(x)=1/(2/3)=3/2$. Finally, substituting this back into our expression for $log_{y}(x^{4})$: $log_{y}(x^{4})=4*(3/2)=6$. Therefore, the value of $log_{y}(x^{4})$ is $6$.
UnKE To find the value of $log_{y}(x^{4})$, we can use the property of logarithms that states: $log_{b}(a)=1/log_{a}(b)$. Given that $log_{x}(y^{3})=2$, we can rewrite this as $y^{3}=x^{2}$. Now, we can use the property to find $log_{y}(x^{4})$: $log_{y}(x^{4})=1/log_{x}(y^{3})$ $=1/(1/x^{2})$ $=x^{2}$. So, the value of $log_{y}(x^{4})$ is $x^{2}$. Therefore, the correct answer is $x^{2}$.

### v3

@@ 7 Conclusion & Limitations
Ackonwledgments
This research is supported by the National Science and Technology Major Project (2023ZD0121102) and the National Natural Science Foundation of China (U24B20180, 62121002). This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ A.1 Datasets
UnKEBench (Deng et al., 2025) constructs a dataset containing 1,000 counterfactual unstructured texts, where knowledge is presented in an unstructured and relatively lengthy form, going beyond simple knowledge triplets or linear fact chains. These texts originate from ConflictQA (Xie et al., 2023), a benchmark specifically designed to distinguish LLMs’ parameter memory from anti-memory. This approach is crucial for preventing the model from merging knowledge obtained during pretraining with knowledge acquired during the editing process. Moreover, it addresses the key challenge of determining whether the model learns target knowledge during training or editing, ensuring a clear boundary between pretraining knowledge and edited knowledge.
@@ A.2 Evaluation Metrics
[…]
@@
[…]
@@
[…]
@@ A.3 Baseline Methods
  • AlphaEdit (Fang et al., 2025) is a method designed to mitigate interference in LLM lifelong knowledge editing. It introduces a null-space projection mechanism that ensures parameter updates preserve previously edited knowledge while incorporating new updates. AlphaEdit has demonstrated state-of-the-art (SOTA) performance across multiple evaluation metrics while maintaining strong transferability. We follow the original paper’s hyperparameter configuration in our implementation.
  • UnKE (Deng et al., 2025) improves knowledge editing by refining both the layer and token dimensions. In the layer dimension, it replaces local key-value storage with a non-local block-based mechanism, enhancing the representation capability of key-value pairs while integrating attention-layer knowledge. In the token dimension, it replaces ”term-driven optimization” with ”cause-driven optimization,” which directly edits the final token while preserving contextual coherence. This eliminates the need for explicit term localization and prevents context loss.
@@ A.4 Implementation Details
Our AnyEdit and AnyEdit* primarily follow the baseline configurations of MEMIT and UnKE, while other baselines adhere to their original implementation settings.
@@ B.1 Locate-Then-Edit Paradigm
Recent studies in LLM behavior analysis (Zhang et al., 2025; Liu et al., 2024b) have explored various aspects of model manipulation, motivating our approach to structured knowledge editing. Following prior works on model editing, the detailed descriptions of specific methods in this section are based on MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025) and ECE (Zhang et al., ). We adhere to their formulations and methodological explanations to ensure consistency and clarity in presenting these approaches.
@@
Methods such as ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), and AlphaEdit (Fang et al., 2025) focus on triplets $(s,r,o)$, selecting the last token of the subject $s$ as the target token. In contrast, UnKE (Deng et al., 2025) extends to unstructured text, using the last token of $x_{e}$ as the target.
@@ C.1.2 Case 2
MEMIT To find the value of $log_{y}(x^{4})$, we can use the change of base formula and the given information. Given: $log_{x}(y^{3})=2$. Using the change of base formula, we can rewrite this as: $log_{x}(y^{3})=(log_{y}(y^{3}))/(log_{y}(x))$. Since $log_{y}(y^{3})=3$ (because $y^{3}$ is the argument raised to the power of 3), we have: $2=3/(log_{y}(x))$. Now, we can solve for $log_{y}(x)$: $log_{y}(x)=3/2$. Next, we want to find $log_{y}(x^{4})$. Using the power rule of logarithms, we can write: $log_{y}(x^{4})=4*log_{y}(x)$. Substituting the value we found for $log_{y}(x)$: $log_{y}(x^{4})=4*(3/2)$. Simplifying: $log_{y}(x^{4})=6$. Therefore, the value of $log_{y}(x^{4})$ is 66.
AlphaEdit To find the value of $log_{y}(x^{4})$, we can use the change of base formula and the given information: $log_{x}(y^{3})=2$. Using the change of base formula, we can rewrite this as: $y^{3}=x^{2}$. Now, we want to find $log_{y}(x^{4})$. Using the properties of logarithms, we can rewrite this as: $log_{y}(x^{4})=4*log_{y}(x)$. We can express $log_{y}(x)$ in terms of $log_{x}(y)$ using the change of base formula: $log_{y}(x)=1/log_{x}(y)$. Since we know that $log_{x}(y^{3})=2$, we can write: $log_{x}(y^{3})=3*log_{x}(y)=2$. Therefore: $log_{x}(y)=2/3$. Substituting this into our expression for $log_{y}(x)$: $log_{y}(x)=1/(2/3)=3/2$. Finally, substituting this back into our expression for $log_{y}(x^{4})$: $log_{y}(x^{4})=4*(3/2)=6$. Therefore, the value of $log_{y}(x^{4})$ is 66.
UnKE To find the value of $log_{y}(x^{4})$, we can use the property of logarithms that states: $log_{b}(a)=1/log_{a}(b)$. Given that $log_{x}(y^{3})=2$, we can rewrite this as $y^{3}=x^{2}$. Now, we can use the property to find $log_{y}(x^{4})$: $log_{y}(x^{4})=1/log_{x}(y^{3})$ $1/(1/x^{2}) =$$x^{2}$. So, the value of $log_{y}(x^{4})$ is $x^{2}$. Therefore, the correct answer is $x^{2}$.